solve for x
1/4x-2=1/2x-3

Answer by jaimehot(1)

1/4x-2=1/2x-3
first add 3 to both sides of the equation
1/4x-2 + 3 =1/2x-3 + 3
you get
1/4x + 1 = 1/2x
now subtract 1/4x to both sides of the equation
1/4x - 1/4x + 1 = 1/2x - 1/4x
or
1 = 1/2x - 1/4x
but we know that
1/2x = 2/4x
substitute in previous equqtion and you have
1 = 2/4x -1/4x
or
1 = 1/4x
multiply both sides times 4 and
4 = x
so
x=4